there are 25 applications for four software tester positions
1
Expert's answer
2018-04-30T08:14:08-0400
If there are 25 applications for four software tester positions, then the number of possible combinations is:
C(25;4) = n!/(m!*(n - m)!) = 25!/(4!*(25 - 4)!) = 22*23*24*25/(1*2*3*4) = 22*23*25 = 12650.